OpenStudy (anonymous):

If the mass per unit area of a surface is given by p=xy , find the mass int(int_S(xy)dS if S is the part of the cylinder x^2+y^2=9 which is in the first octant and contained within the cylinder x^2+y^2=1 need some help setting this up, got it wrong multiple times

OpenStudy (anonymous):

The cylinder x^2+y^2=9 has no intersection with cylinder x^2+y^2=1 so this makes no sense to me "S is the part of the cylinder x^2+y^2=9" and "contained within the cylinder x^2+y^2=1"

OpenStudy (anonymous):

There is something wrong with your cylinder equations or you need a z bound. Something about the problem isn't right.

OpenStudy (anonymous):

I hope; you have committed mistake while typing. I think your range will be x^2+y^2=9 & x^2+z^2=1; then ans. will be 3.